# Progressive Autoregressive Video Diffusion Models

### v1

@@ Abstract
Current frontier video diffusion models have demonstrated remarkable results at generating high-quality videos. However, they can only generate short video clips, normally around 10 seconds or 240 frames, due to computation limitations during training. In this work, we show that existing models can be naturally extended to autoregressive video diffusion models without changing the architectures. Our key idea is to assign the latent frames with progressively increasing noise levels rather than a single noise level, which allows for fine-grained condition among the latents and large overlaps between the attention windows. Such progressive video denoising allows our models to autoregressively generate video frames without quality degradation or abrupt scene changes. We present state-of-the-art results on long video generation at 1 minute (1440 frames at 24 FPS). Videos from this paper are available at https://desaixie.github.io/pa-vdm/.
@@ 1 Introduction
Video diffusion models have recently demonstrated remarkable success in generating high-quality video content across a variety of applications (Meta, 2024; Blattmann et al., 2023a; Guo et al., 2023; Bar-Tal et al., 2024; Ho et al., 2022a; Girdhar et al., 2023). These models are capable of synthesizing realistic video sequences that are increasingly indistinguishable from real-world footage. However, despite their impressive results, current video diffusion models are constrained by a significant limitation: they can only generate videos of relatively short duration, typically up to about 10 seconds or 240 frames. This temporal restriction leads to challenges for broader applications that require longer, more continuous video outputs, highlighting the need for further research and innovation to extend the capabilities of these models.
Several approaches (Ho et al., 2022b; Henschel et al., 2024; Blattmann et al., 2023a; Zheng et al., 2024; Gao et al., 2024) have been proposed to address the challenge of generating longer videos with video diffusion models. They autoregressively apply video diffusion models to generate short video clips in a windowed fashion, where each subsequent clip conditions on the final frames of the previous one. One solution (Zheng et al., 2024; Gao et al., 2024) directly puts the conditioning frames into the input frames, replacing the noisy frames. Another solution (Ho et al., 2022b) additionally adds the same level of noise to the conditioning frames as the noisy frames. Due to the large gap between the noise levels of the condition frames and the noisy frames, iteratively applying the first solution sometimes produces videos with numerous abrupt scene changes, which is similar to generating short video clips independently. The second solution is more effective at producing smooth pixel-level transitions between generated clips, since the condition frames and the noisy frames have the same noise levels during denoising. However, it struggles to preserve secondary attributes, e.g., velocity and acceleration, accurately, leading to generating unnatural and inconsistent movement in long videos. In addition, repeatedly applying the second solution amplifies these issues, increasing inaccuracies in motion dynamics and accumulating errors that causes the video to eventually diverge.
We propose autoregressive video diffusion models that denoise video frames in a progressive manner. It can generate long videos with smooth and continuous video frames and motions with no quality degradation. The core innovation of our method lies in the denoising process: instead of applying a single noise level across all frames used in traditional video generation (or extension) diffusion models, we progressively increase the noise levels across frames during denoising. Such progressive noise levels allows autoregressive video denoising with large overlaps between the attention window, as illustrated in Fig. 1. This approach improves the temporal transitions in extended frames, resulting in more natural motion and better consistency. Our method can be easily implemented by changing the noise scheduling and finetuning pre-trained video diffusion models, either UNet-based (Ronneberger et al., 2015; Ho et al., 2022b) or Diffusion Transformer (DiT)-based (Peebles & Xie, 2023; Ma et al., 2024; Brooks et al., 2024) backbone, without changing the original model architecture. Our progressive autoregressive inference procedure can work training-free, if the model has been trained on vary noise levels, e.g. the masked pre-training in (Zheng et al., 2024). Our method can generate videos up to one minute in length (1440 frames) without noticeable degradation in quality. Moreover, the additional computational cost at inference time is minimal comparing to previous works (Wang et al., 2023; Qiu et al., 2024; Henschel et al., 2024) that generates overlapped clips, making this approach more efficient for practical use in the long video generation.
We compare our method to the baselines on long video generation at 60-second or 1440 frames. Our quantitative results demonstrate that our method has overall the best frame quality, the best motion dynamics, the best aesthetic and the least scene changes. It is able to maintain these metrics over the entire 60-second duration. Qualitatively, our method is also shown to be preserve the frame and motion quality over time, while the baselines performance drops. By applying our method to two base models and demonstrating significantly improved performance compared to their respective baselines, we confirm its universal applicability to existing video diffusion models. We provide more video results at our website https://desaixie.github.io/pa-vdm/. In sum, our progressive autoregressive video diffusion models achieves state-of-the-art results on long video generation. To facilitate future research, we release our training and inference code based on Open-Sora (Zheng et al., 2024) at https://github.com/desaixie/pa_vdm.
(a) replacement
(b) ours
Figure 1: Comparison of autoregressively applying video diffusion models with replacement methods (left) vs. our progressive autoregressive video diffusion models (right). Our method allows for more fine-grained condition among the latent frames, where the later frames with more uncertainty can follow the pattern of the earlier, more certain frames, and larger overlap between the attention windows without extra computation cost.
@@ 2.1 Video Diffusion Models
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) are generative models that learn to generate samples from a data distribution $q({\mathbf{x}}^{0})$ through an iterative denoising process. During training, data samples are first corrupted using the forward diffusion process $q({\mathbf{x}}^{t}|{\mathbf{x}}^{0})$
@@
where $t\in[0,T)$ is the noise level or diffusion timestep, $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},{\bm{I}})$ is the noise, and $\bm{\alpha}^{1:T}$ is the variance schedule. With those noisy data samples ${\mathbf{x}}^{t}$, diffusion models are trained to fit to the data distribution $q({\mathbf{x}}^{0})$ by maximizing the variational lower bound (Kingma & Welling, 2013) of the log likelihood of ${\mathbf{x}}^{0}$, which can be simplified into a mean squared error loss (Ho et al., 2020)
@@
At sampling time, we consider the sampling noise level schedule $\bm{\tau}=\{\tau_{1},\tau_{2},...,\tau_{S}\}$ , which is an increasing subset of $t\in[0,T)$ of length $S$ (Song et al., 2021). Starting from ${\mathbf{x}}^{\tau_{S}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}),\tau_{S}=T$, the reverse denoising process is autoregressively applied as
@@
where $\hat{{\mathbf{x}}}^{0}=\displaystyle f_{\theta}({\mathbf{x}}^{t},t)$ is the ${\mathbf{x}}^{0}$ predicted by the model and $f_{\theta}({\mathbf{x}}^{t},t)$ is the DDIM (Song et al., 2021) reverse process equation, which we omit for simplicity. This gives us a sequence of samples ${\mathbf{x}}^{T},{\mathbf{x}}^{\tau_{S-1}},\ldots,{\mathbf{x}}^{\tau_{1}},{\mathbf{x}}^{0}$, and the last sample ${\mathbf{x}}^{0}$ is the clean output result.
Video diffusion models (Ho et al., 2022b; Blattmann et al., 2023b) are diffusion models that consider latent representations of video data, consisting of $F$ latent frames ${\mathbf{x}}_{0:F-1}=\{{\textnormal{x}}_{0},{\textnormal{x}}_{1},\ldots,{\textnormal{x}}_{F-1}\}$. The same forward diffusion process, the reverse denoising process, and loss (eqs. 1, 2, 3 and 4) can be applied to model these video data by treating all the latents as one entity, ignoring the correlation among the latents. Recent video diffusion models (Meta, 2024; Zheng et al., 2024) have employed various diffusion model variants (Liu et al., 2022; 2023; Lipman et al., 2023) to improve training and inference efficiency as well as output quality. Nevertheless, our method is compatible with any diffusion model variant as long as the model corrupts the data ${\mathbf{x}}^{t}$ at the same noise levels $t$.
2.2 Long Video Generation via Replacement
Video diffusion models can only generate short video clips, because they are only trained on videos with a limited length $F$ due to GPU memory limit. When adapted to generating $L>F$ latent frames at sampling time, their generation quality substantially degrades (Qiu et al., 2024). The straightforward solution is to autoregressively apply video diffusion models, generating each video clip while conditioning on the previous clip. In this paper, we refer to the $F$ latents that the video diffusion model processes as the attention window.
Given $E<F$ clean latents ${\mathbf{x}}_{0:E}^{0}$ as condition, there are two methods for autoregressively applying video diffusion models. Zheng et al. (2024); Gao et al. (2024); Blattmann et al. (2023a) place the clean condition latents $\bar{{\mathbf{x}}}_{0:E-1}^{0}$ directly at the front of the attention window, directly replacing the sampled latents ${\mathbf{x}}_{0:E-1}^{\tau_{i}}$ at each denoising step
@@ 2.2 Long Video Generation via Replacement
Song et al. (2020b); Ho et al. (2022b) additionally add noise to the conditioning latents
@@
where $\bar{{\mathbf{x}}}_{0:E-1}^{\tau_{i}}$ are the condition latents $\bar{{\mathbf{x}}}_{0:E-1}^{0}$ noised via the forward process (eqs. 1 and 2). This maintains the same noise level distribution and training objective as regular video diffusion models. We will refer to this method as the replacement-with-noise method.
@@
Figure 2: Comparison of noise levels of ours vs. the replacement without noise method.
@@ 3 Progressive Autoregressive Video Diffusion Models
Existing video diffusion models (Meta, 2024; Blattmann et al., 2023a; Guo et al., 2023; Bar-Tal et al., 2024) can only generate short video clips up to a limited length (e.g. 10 seconds or 240 frames), due to GPU memory constraints during training. We show that, without changing the architectures, they can be naturally extended to progressive autoregressive video diffusion models that can generate long videos without quality degradation. We achieve this by proposing a per-frame noise schedule, which is inspired by Chen et al. (2024). During training, we finetune pre-trained video diffusion models to adapt to such noise schedule; during sampling, our models adopt such noise schedule and can thus autoregressively generate video frames.
3.1 Progressive Video Denoising
Inspired by (Chen et al., 2024), we assign progressively increasing noise levels $\bm{\tau}_{0:F-1}$ to the $F$ latent frames in the attention window. These noise levels can be obtained from the sampling noise level schedule $\bm{\tau}=\{\tau_{0},\tau_{1},...,\tau_{S}\}$ (Song et al., 2021), where $0=\tau_{0}<\ldots<\tau_{i-1}<\tau_{i}<\ldots<\tau_{S}=T$, depending on $F$ and $S$. In this work, we consider the linear sampling schedule
$\displaystyle\bm{\tau}^{\prime}_{0:S}=\left\{\tau^{\prime}_{0},\tau^{\prime}_{1},\tau^{\prime}_{2},\ldots,\tau^{\prime}_{S-1},\tau^{\prime}_{S}\right\}=\left\{0,\frac{T}{S},\frac{2T}{S},\ldots,\frac{(S-1)T}{S},T\right\}$ (7)
For simplicity, we set $F=S$, so that our latents’ noise levels $\bm{\tau}_{0:F-1}$ can take the values of either $\bm{\tau}^{\prime}_{0:S-1}$ or $\bm{\tau}^{\prime}_{1:S}$. We denoise the latent frames of a video with the progressively increasing noise levels
$\displaystyle p_{\theta}\left({\textnormal{x}}_{0}^{\tau^{\prime}_{0}},{\textnormal{x}}_{1}^{\tau^{\prime}_{1}},...,{\textnormal{x}}_{F-2}^{\tau^{\prime}_{S-2}},{\textnormal{x}}_{F-1}^{\tau^{\prime}_{S-1}}\middle|{\textnormal{x}}_{0}^{\tau^{\prime}_{1}},{\textnormal{x}}_{1}^{\tau^{\prime}_{2}},...,{\textnormal{x}}_{F-2}^{\tau^{\prime}_{S-1}},{\textnormal{x}}_{F-1}^{\tau^{\prime}_{S}}\right)$ (8)
where the input frames and output frames have noise levels $\bm{\tau}^{\prime}_{1:S}$ and $\bm{\tau}^{\prime}_{0:S-1}$ respectively.
With our progressive noise levels, which is a subset of independent noise levels ${\mathbf{t}}_{0:F-1}$ for the frames ${\mathbf{x}}_{0:F-1}$, one concern is that whether this formulation is still compatible with existing video diffusion models. Conceptually, we are effectively using a single set of model parameters $\theta$ to jointly model $F$ latents $q({\textnormal{x}}^{0}_{0}),q({\textnormal{x}}^{0}_{1}),\ldots,q({\textnormal{x}}^{0}_{F-1})$, where each latent has a single noise level $t_{f}$ like regular diffusion models. Thus, we can directly apply the forward process, reverse process, and training loss (eqs. 1, 2, 3 and 4) of existing video diffusion models (Ho et al., 2020; Song et al., 2021; Liu et al., 2022; 2023). This means that we can obtain our progressive autoregressive video diffusion models from pre-trained video diffusion models, which demands immense computation resources.
Intuitively, the benefit of our progressive video denoising process is that it gradually establishes correlation among consecutive latent frames. Given some existing video latents as conditioning, it is challenging for video diffusion models to produce temporally consistent extensions latents from newly sampled noisy latents (Qiu et al., 2024). In contrast to the replacement-with-noise method (Blattmann et al., 2023a; Ho et al., 2022b) where the latents are denoised together at the same noise level, our progressive video denoising encourages the later latents with higher uncertainty to follow the patterns of the earlier and more certain latents, facilitating modeling a smoother temporal transition and better preserving motion velocity. Compared to the replacement-without-noise method where there is a large noise level gap between the clean condition latents $\bar{{\mathbf{x}}}_{0:E-1}^{0}$ and the noisy frames ${\mathbf{x}}_{E:F-1}^{\tau_{i}}$, our method provides more fine-grained condition, where difference of noise levels between neighboring latents is only $\frac{T}{S}$, as illustrated in eqs. 7 and 2. Additionally, both replacement methods (Ho et al., 2022b; Blattmann et al., 2023a; Zheng et al., 2024) need to balance between the benefit of having larger overlap between two subsequent attention windows and the extra computation cost when setting the number of condition frames $E<F$. In contrast, our method can have the maximum attention window overlap of $F-1$ latents while not requiring any extra computation (eq. 8).
Another benefit of our method is the more efficient modeling of long videos of $L$ latent frames using video diffusion models with a limited attention window length $F<L$. During training, the model needs to fit any subset of $F$ latents out of the $L$ latents. When using the two replacement methods (Ho et al., 2022b; Blattmann et al., 2023a; Zheng et al., 2024), the model needs to train on the full range of noise levels $t\in[0,T)$ for each data sample of $F$ latents. With our method, the neighboring training data are efficiently modeled as exactly one sampling step apart from each other, so that we only need to train on the noise levels between $\bm{\tau}^{\prime}_{0:S-1}$ and $\bm{\tau}^{\prime}_{1:S}$.
3.2 Autoregressive Generation
Notice that the noise levels of the input and output latents in eqs. 7 and 8, $\bm{\tau}^{\prime}_{0:S-1}$ and $\bm{\tau}^{\prime}_{1:S}$, only differ by $\tau^{\prime}_{0}=0$ and $\tau^{\prime}_{S}=T$. We can simply transition the output latents into the correct input noise levels by saving and removing the clean latent ${\textnormal{x}}_{0}^{0}$ at the front, shifting the latent sequence forward once, and appending a new noisy latent ${\textnormal{x}}_{F-1}^{T}\sim\mathcal{N}(\mathbf{0},{\bm{I}})$ at back. Algs. 1 and 1 details this autoregerssive sampling procedure.
@@ 3.2 Autoregressive Generation
2:$\bm{\tau}^{\prime}_{0:S}=\left\{\tau^{\prime}_{0},\tau^{\prime}_{1},\ldots,\tau^{\prime}_{S}\right\}=\left\{0,\frac{T}{S},\ldots,T\right\}$triangleright eq. 7, progressive noise levels
@@
4:${\mathbf{x}}^{\bm{\tau}^{\prime}_{1:S}}_{0:F-1}=\sqrt{\alpha^{\bm{\tau}^{\prime}_{1:S}}}{\mathbf{x}}^{0}_{0:F-1}+\sqrt{1-\alpha^{\bm{\tau}^{\prime}_{1:S}}}\bm{\epsilon}$triangleright eq. 2, add noise
@@
6:     ${\mathbf{x}}^{\bm{\tau}^{\prime}_{0:S-1}}_{0:F-1}=\left\{{\textnormal{x}}_{0}^{0},{\textnormal{x}}_{1}^{\tau^{\prime}_{1}},\ldots,{\textnormal{x}}_{F-1}^{\tau^{\prime}_{S-1}}\right\}\sim p_{\theta}\left({\mathbf{x}}^{\bm{\tau}^{\prime}_{0:S-1}}_{0:F-1}\middle|{\mathbf{x}}^{\bm{\tau}^{\prime}_{1:S}}_{0:F-1}\right)$triangleright eq. 8, one sampling step
7:     ${\textnormal{x}}_{F-1}^{T}\sim\mathcal{N}(\mathbf{0},{\bm{I}})$triangleright Sample a new noisy latent
8:     Append ${\textnormal{x}}_{0}^{0}$ to the list of clean latents
9:     ${\mathbf{x}}^{\bm{\tau}^{\prime}_{0:S}}_{0:F-1}=\left\{{\textnormal{x}}_{1}^{\tau^{\prime}_{1}},\ldots,{\textnormal{x}}_{F-2}^{\tau^{\prime}_{S-1}},{\textnormal{x}}_{F-1}^{T}\right\}$triangleright Remove ${\textnormal{x}}_{0}^{0}$, shift latents by $-1$, and append ${\textnormal{x}}_{F-1}^{T}$

### v2

@@ Abstract
Current frontier video diffusion models have demonstrated remarkable results at generating high-quality videos. However, they can only generate short video clips, normally around 10 seconds or 240 frames, due to computation limitations during training. Existing methods naively achieve autoregressive long video generation by directly placing the ending of the previous clip at the front of the attention window as conditioning, which leads to abrupt scene changes, unnatural motion, and error accumulation. In this work, we introduce a more natural formulation of autoregressive long video generation by revisiting the noise level assumption in video diffusion models. Our key idea is to 1. assign the frames with per-frame, progressively increasing noise levels rather than a single noise level and 2. denoise and shift the frames in small intervals rather than all at once. This allows for smoother attention correspondence among frames with adjacent noise levels, larger overlaps between the attention windows, and better propagation of information from the earlier to the later frames. Video diffusion models equipped with our progressive noise schedule can autoregressively generate long videos with much improved fidelity compared to the baselines and minimal quality degradation over time. We present the first results on text-conditioned 60-second (1440 frames) long video generation at a quality close to frontier models. Code and video results are available at https://desaixie.github.io/pa-vdm/.
†This work is done while Desai is an intern at Adobe Research.
@@ 1 Introduction
(a) the replacement method
(b) our PA-VDM
Figure 1: Comparison of autoregressive long video generation methods. Top: the replacement method, which replaces the front of noisy latent frames with the ending of previous clip as condition and denoising all the frames at once. Bottom: our PA-VDM, which applies progressive noise levels and denoises and shifts the frames in small intervals. The final long video consists of autoregressively generated clean frames. $\oplus$ denotes concatenation. The noise level $t$ for each frame is illustrated by the solid color of the frame, where darker colors are closer to $1$ and lighter colors are closer to 00.
@@
Frontier video diffusion models [3, 34, 9, 55, 31, 29, 22, 24, 21, 51, 39] have recently demonstrated remarkable success in generating high-quality video contents by scaling up transformer-based [32, 48] architectures. However, they can only generate videos of relatively short duration, typically up to about 10 seconds or 240 frames, due to the demanding computation cost of long-sequence training. This temporal restriction leads to challenges for broader applications that require longer, more continuous video outputs.
Several approaches [15, 12, 1, 58, 7] have been proposed to autoregressively apply video diffusion models for long video generation; they generate short video clips in a windowed fashion, where each subsequent clip conditions on the final frames of the previous one. One solution [58, 7] directly places the conditioning frames into the input frames, replacing the noisy frames. Another solution [43, 15] additionally adds the same level of noise to the conditioning frames as the noisy frames. This naive way of conditioning suffers from various flaws, including temporal inconsistency, abrupt scene changes, unnatural motion dynamics, and accumulated errors that lead to divergence.
In this work, we propose Progressive Autoregressive Video Diffusion Models (PA-VDM) for high-quality long video generation. The core innovation of our method lies in the denoising process: instead of applying a single noise level across all frames used in traditional video diffusion models [15, 2], we apply progressively increasing noise levels across the frames; correspondingly, we denoise and shift the frames in small intervals, instead of denoising and shift them all at once. We illustrate our method in Fig. 1. Such progressive noise levels and autoregressive video denoising benefit from larger overlaps between subsequent attention windows, smoother attention correspondence among frames with adjacent noise levels, and better propagation of information from the earlier to the later frames. When applying our variable length progressive noise schedule, our models can start or end the autoregressive generation at arbitrary video lengths. Our chunked frames and overlapped conditioning techniques prevent divergent results and chunk-to-chunk discontinuity. Together, our method can autoregressively generate long videos while maintaining the initial quality over time.
PA-VDM provides a range of benefits for the video generation community. It can be easily implemented by changing the noise scheduling and finetuning pre-trained video diffusion models without changing the original model architecture; this allows our method to be easily reproduced and combined with orthogonal methods, such as external memory modules [12] and multiple text prompts [59, 11]. While we choose to demonstrate PA-VDM on Diffusion Transformer (DiT)-based [32, 30, 3] models, PA-VDM is model agnostic and can be extended to UNet-based [37, 15] models. As shown in Sec. 4.2, our method can work training-free, if the model has been trained on varied noise levels [58]. Moreover, the additional inference computational cost of PA-VDM is minimal without sacrificing any generation quality, as opposed to previous works [49, 36, 12] that need to trade off quality for efficiency, making this approach more efficient for practical use in long video generation.
We compare our method to the baselines on a text-conditioned 60-second (1440 frames) long video generation benchmark consisting of 40 real videos and their captions. Our quantitative results demonstrate that our results have overall the best quality across various dimensions and are the best at maintaining these metrics over the entire 60-second duration. Qualitatively, our method substantially outperforms the baselines in terms of temporal consistency, motion dynamics, and maintaining quality over time. In human evaluation, our models are also favored over various baseline models. Our ablation studies demonstrate the effectiveness of our chunked frames and overlapped conditioning techniques at preventing cumulative error and temporal jittering, respectively. By applying our method to two base models and outperforming their respective baselines, we confirm its universal applicability to existing video diffusion models. We encourage readers to check out our project webpage for video results qualitatively comparing ours and the baselines. To facilitate future research, we also release our code based on Open-Sora [58].
We summarize our contribution as follows:
  1. 1.
    We propose a progressive noise level schedule, an autoregressive video denoising algorithm, and the chunked frames and overlapped conditioning techniques. Together, these enable high-quality long video generation building upon pre-trained video diffusion models.
  2. 2.
    We are the first to achieve 60-second long video generation with quality that are close to frontier models, when compared at the same resolution. On our 60-second long video generation benchmark, we achieve superior VBench and FVD scores, majority preference in human evaluations, and strong qualitative results. This marks a significant step forward in generating longer videos, a dimension that has not been explored by recent frontier video diffusion models [34, 31, 29, 9, 55, 22].
  3. 3.
    Our method benefits the video generation research community in many ways, including easy implementation and reproduction, training-free application, minimal additional inference cost, and universal applicability on video diffusion models.
@@ 2.1 Video Diffusion Models
Diffusion models [40, 13] are generative models that learn to generate samples from a data distribution $q({\mathbf{x}}^{0})$ through an iterative denoising process. During training, data samples are first corrupted using the forward diffusion process $q({\mathbf{x}}^{t}|{\mathbf{x}}^{0})$
@@
where $t\in[0,T)$ is the noise level or diffusion timestep, $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},{\bm{I}})$ is the noise, and $\bm{\alpha}^{1:T}$ is the variance schedule. With those noisy data samples ${\mathbf{x}}^{t}$, diffusion models are trained to fit to the data distribution $q({\mathbf{x}}^{0})$ by maximizing the variational lower bound [20] of the log likelihood of ${\mathbf{x}}^{0}$, which can be simplified into a mean squared error loss [13]
@@
At sampling time, we consider the sampling noise level schedule $\bm{\tau}=\{\tau_{0},\tau_{1},...,\tau_{S}\}$ , which is a monotonically increasing subset of $t\in[0,T)$ of length $S+1$ [42]. Starting from ${\mathbf{x}}^{\tau_{S}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}),\tau_{S}=T$, the reverse denoising process is iteratively applied as
@@
where $\hat{{\mathbf{x}}}^{0}=\displaystyle f_{\theta}({\mathbf{x}}^{t},t)$ is the ${\mathbf{x}}^{0}$ predicted by the model and $f_{\theta}({\mathbf{x}}^{t},t)$ is the DDIM [42] reverse process equation, which we omit for simplicity. This gives us a sequence of samples ${\mathbf{x}}^{T},{\mathbf{x}}^{\tau_{S-1}},\ldots,{\mathbf{x}}^{\tau_{1}},{\mathbf{x}}^{0}$, and the last sample ${\mathbf{x}}^{0}$ is the clean output result.
Latent video diffusion models [15, 2] are diffusion models that models latent representations of video data, consisting of $F$ latent frames ${\mathbf{x}}_{0:F-1}=\{{\textnormal{x}}_{0},{\textnormal{x}}_{1},\ldots,{\textnormal{x}}_{F-1}\}$. The video latent frames are usually spatially and temporally [57] compressed through a VAE [20]. For simplicity, we refer to latent video diffusion models as video diffusion models and latent frames as frames. The same forward process, reverse process, and loss (Eqs. 1, 2, 3 and 4) can be applied to model these video data by treating all the frames as one entity, ignoring the correlation among the frames. Recent video diffusion models [34, 58] have employed various diffusion model variants [26, 27, 25] to improve training and inference efficiency as well as output quality. Nevertheless, our method is compatible with any diffusion model variant as long as the model corrupts the data ${\mathbf{x}}^{t}$ at the same noise levels $t$.
2.2 Autoregressive Long Video Generation via Replacement
Video diffusion models can only generate short video clips, because they are only trained on videos with a limited length $F$ due to GPU memory limit. When adapted to generating $L>F$ latent frames at sampling time, their generation quality substantially degrades [36]. The straightforward solution is to autoregressively apply video diffusion models, generating each video clip while conditioning on the previous clip. In this paper, we refer to the $F$ frames that the video diffusion model processes as the attention window.
Given $E<F$ clean frames ${\mathbf{x}}_{0:E}^{0}$ as condition, there are two methods for autoregressively applying video diffusion models. [58, 7, 1] place the clean condition frames $\bar{{\mathbf{x}}}_{0:E-1}^{0}$ directly at the front of the attention window, directly replacing the sampled frames ${\mathbf{x}}_{0:E-1}^{\tau_{i}}$ at each denoising step
@@ 2.2 Autoregressive Long Video Generation via Replacement
[43, 15] additionally add noise to the condition frames
@@
where $\bar{{\mathbf{x}}}_{0:E-1}^{\tau_{i}}$ are the condition frames $\bar{{\mathbf{x}}}_{0:E-1}^{0}$ noised via the forward process (Eqs. 1 and 2). This maintains the same noise level distribution and training objective as regular video diffusion models. We will refer to this method as the replacement-with-noise method. Note that [15] proposes reconstruction guidance for the replacement-with-noise method but is not widely adopted.
@@
See Appendix B for a detailed discussion of two parallel works [38, 19] that share a high-level idea similar to our work. Please refer to Appendix C for related works.
Figure 2: Comparison of noise levels of a sequence of video frames when using the replacement without noise method (left) and ours (right).
@@ 3 Progressive Autoregressive Video Diffusion Models
2:$\bm{\tau}_{0:S}=\left\{\tau_{0},\tau_{1},\ldots,\tau_{S}\right\}=\left\{0,\frac{T}{S},\ldots,T\right\}$triangleright Eq. 7, linear sampling noise level schedule
@@
4:${\mathbf{x}}^{\bm{\tau}_{1:S}}_{0:F-1}=\sqrt{\alpha^{\bm{\tau}_{1:S}}}{\mathbf{x}}^{0}_{0:F-1}+\sqrt{1-\alpha^{\bm{\tau}_{1:S}}}\bm{\epsilon}$triangleright Eq. 2, add noise and set to progressive noise levels
@@
6:     ${\mathbf{x}}^{\bm{\tau}_{0:S-1}}_{0:F-1}=\left\{{\textnormal{x}}_{0}^{0},{\textnormal{x}}_{1}^{\tau_{1}},\ldots,{\textnormal{x}}_{F-1}^{\tau_{S-1}}\right\}\sim p_{\theta}\left({\mathbf{x}}^{\bm{\tau}_{0:S-1}}_{0:F-1}\middle|{\mathbf{x}}^{\bm{\tau}_{1:S}}_{0:F-1}\right)$triangleright Eq. 8, one sampling step
7:     ${\textnormal{x}}_{F-1}^{T}\sim\mathcal{N}(\mathbf{0},{\bm{I}})$triangleright Sample a new noisy frame
8:     Append ${\textnormal{x}}_{0}^{0}$ to the list of clean frames
9:     ${\mathbf{x}}^{\bm{\tau}_{0:S}}_{0:F-1}=\left\{{\textnormal{x}}_{1}^{\tau_{1}},\ldots,{\textnormal{x}}_{F-2}^{\tau_{S-1}},{\textnormal{x}}_{F-1}^{T}\right\}$triangleright Remove ${\textnormal{x}}_{0}^{0}$, shift frames forward, and append ${\textnormal{x}}_{F-1}^{T}$